ok i do not have an specific question it's just that i have to take the crct for the second time im in summer school and we take the test next tuesday i was just wondering do you have any worksheets on pythagorean theorem please and thank you
Answer by rfer(16322)

just remember
a^2+b^2=c^2
most of the time they will use what is called triples, which are
3,4,5 or 6,8,10 triangle
That means that the side are 3,4 and the hypotenuse is 5
(2)^2+(6)^2=(x)^2
4+36=40
then take the square root of 40 for your answer =6.32